In January 2016, United Airlines (UAL) had a market capitalization of $20.66 billion, debt of $12.01 billion, and cash of $5.09 billion. United Airlines had revenues of $39.58 billion. Southwest Airlines (LUV) had a market capitalization of $27.49billion, debt of $3.17 billion, cash of $3.06 billion, and revenues of $19.83 billion. Southwest Airlines. b. Compare the enterprise value-to-revenue ratio for United Airlines and Southwest Airlines. c. Which of these comparisons is more meaningful? Explain. a. Compare the market capitalization-to-revenue ratio (also called the price-to-sales ratio) for United Airlines and Southwest Airlines.
Answer #1

(a) United Airlines:

Market Cap = $ 20.66 billion and Revenue = $ 39.58 billion

Market Cap / Revenue = 20.66 / 39.58 ~ 0.522

Southwest Airlines:

Market Cap = $ 27.49 billion and Revenue = $ 19.83 billion

Market Cap / Revenue = 27.49 / 19.83 ~ 1.386

(b) United Airlines:

Enterprise Value (EV) = Market Cap + Debt - Cash = 20.66 + 12.01- 5.09 = $ 27.58 billion

EV / Revenue = 27.58 / 39.58 ~ 0.697

Southwest Airlines:

Enterprise Value (EV) = Market Cap + Debt - Cash = 27.49 + 3.17 - 3.06 = $ 27.6 billion

EV / Revenue = 27.6 / 19.83 ~ 1.392

(c) The EV / Revenue ratio is more meaningful as Revenue encompasses the cash flows coming from both equity and debt (sources of capital) which is consistent with the fact that EV encompasses the value of both equity and debt. The Market Cap / Revenue Ratio is incosistent because the numerator is the price of only the equity held by the firm whereas the denominator encompasses cash flows accruing to both equity and debtholders.
Answer #2

another answer for C

C) The market capitalization  to revenue ratio cannot be meaningfully compared when the firms have different amounts of leverage, as market capitalization measures only the value of the firm's equity. The enterprise value to revenue ratio is therefore more useful when firm's leverage is quite different, as it is here.